# New simple and accurate approximations for quintessence

### v2

@@ 3.2 Tracking quintessence
Next, we introduce a dynamical system that is suitable for deriving approximations for the evolution in the vicinity of $\phi=0$ and as long the potential and $\lambda$ are monotonically decreasing, i.e., when $\lambda(\phi)>0$ and $\Gamma(\phi)>1$, as follows from (40) and (48). The system of equations are based on the following new variables: [^14] The variables $\tilde{\phi}$ and $v$ are slightly different than the variables $\bar{\varphi}$ and $v$ in [32], which were adapted to a global state space setting, while the present ones are designed for fast convergence for the tracker orbit and the explicit appearance of $\Gamma$ in the equations and results.
@@
The dynamical system (52) admits $\tilde{\phi}=0$ (and thereby $\phi=0$) as an invariant boundary subset. As shown in [32], in regularizing variables tracking quintessence is associated with a tracker orbit that originates from a tracker fixed point, $\mathrm{T}$, on the $\tilde{\phi}=0$ boundary, which in the present variables is given by
@@
Notably $\mathrm{T}$ is a hyperbolic saddle with the ‘tracker orbit’ as its unstable manifold and $\tilde{\phi}=0$ being its 2D stable manifold, which results in that the tracker orbit acts as an ‘attractor orbit’, see [32] for details.
To find the parameters $w_{\infty}$, $\gamma$ and $\beta$ we first note that the tracker fixed point value for $u=u_{\mathrm{T}}$ gives the initial asymptotic tracker value for $w_{\phi}$ since
@@
$\Gamma(\tilde{\phi})\approx\Gamma^{(0)}+\Gamma^{(1)}\tilde{\phi}+\frac{1}{2}\Gamma^{(2)}\tilde{\phi}^{2}+\dots,\qquad\Gamma^{(n)}:=\left.\frac{d^{n}\Gamma}{d\tilde{\phi}^{n}}\right|_{\tilde{\phi}=0},$ (56)
@@
where we in the expression for $k^{(2)}$ assume that $k\neq 0$. Note that the special class of models with $k=0$, i.e. $\Gamma^{(1)}=\frac{3}{2}w_{\infty}$, needs special treatment and that it include models with $w_{\phi}=\mathrm{constant}=w_{\infty}$, see [32] for details. [^15] Sahni and Starobinsky (2000) [47], equation (121), and Urena-Lopez and Matos (2000) [48], investigated models with the potential $V=V_{0}\sinh^{-p}(\nu\phi)$ and showed that these models admit a special solution (which turns out to be the tracker orbit) for which $w_{\phi}$ is constant, i.e., they yield $w$CDM spacetimes with $-1<w_{\phi}=w_{\infty}=w<0$.
@@ 4 Numerical comparisons
In this section we calculate the numerical relative errors of the approximations, i.e.,
@@
of our approximations for thawing and tracking quintessence, but also for approximations found in the literature, described in appendix A. For brevity we will consider the relative errors of $w_{\phi}(z)$, $\Omega_{\mathrm{m}}(z)$ and $E(z)$.
Before doing so, let us write our approximations for $w_{\mathrm{DE}}$, $\Omega_{\mathrm{DE}}$, from which $\Omega_{\mathrm{m}}$ follows since $\Omega_{\mathrm{m}}=1-\Omega_{\mathrm{DE}}$, and $E_{\mathrm{DE}}^{2}$ in the redshift $z$. The approximations based on $[0/1]_{w_{\mathrm{DE}}}(T)$ and only involving $\gamma$ are given by (as follows from (26a), (30a) and (30b)):
$w_{\mathrm{DE}}$ $\approx w_{\infty}\left(1-\frac{(\gamma-1)\Omega_{\mathrm{DE}0}}{(\gamma-1)\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})(1+z)^{-3w_{\infty}}}\right),$ (59a)
$\Omega_{\mathrm{DE}}$ $\approx\frac{\Omega_{\mathrm{DE},0}}{\gamma\Omega_{\mathrm{DE},0}+(1-\gamma\Omega_{\mathrm{DE},0})(1+z)^{-3w_{\infty}}},$ (59b)
$E_{\mathrm{DE}}^{2}$ $\approx\Omega_{\mathrm{m}0}(1+z)^{3}\left(1+\frac{\Omega_{\mathrm{DE}0}}{(\gamma-1)\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})(1+z)^{-3w_{\infty}}}\right),$ (59c)
while the approximations obtained from $[1/1]_{w_{\mathrm{DE}}}(T)$, involving both $\gamma$ and $\beta$, are as follows:
$w_{\mathrm{DE}}$ $\approx w_{\infty}\left(1-\frac{(\gamma-1)\Omega_{\mathrm{DE}0}}{\beta\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})(1+z)^{-3w_{\infty}}}\right),$ (60a)
$\Omega_{\mathrm{DE}}$ $\approx\frac{\Omega_{\mathrm{DE}0}}{\Omega_{\mathrm{m}0}(1+z)^{-3w_{\infty}}\left(\frac{1+\beta T}{1+\beta T_{0}}\right)^{\frac{\gamma-1}{\beta}}+\Omega_{\mathrm{DE}0}},$ (60b)
$E_{\mathrm{DE}}^{2}$ $\approx(1+z)^{3}\left[\Omega_{\mathrm{m}0}+\Omega_{\mathrm{DE}0}(1+z)^{3w_{\infty}}\left(\frac{1+\beta T_{0}}{1+\beta T}\right)^{\frac{\gamma-1}{\beta}}\right],$ (60c)
where
$T=T_{0}(1+z)^{3w_{\infty}},\qquad T_{0}=\frac{\Omega_{\mathrm{DE}0}}{1-\gamma\Omega_{\mathrm{DE}0}}.$ (60d)
The parameters $\gamma$ and $\beta$ for thawing quintessence were given in equations (46) and (47), while those for tracking quintessence are found in equations (57), where we recall that in the tracking quintessence case $w_{\infty}=-2/(2+p)$ where $p=\lim_{\phi\rightarrow 0}(\phi\lambda)$. We cannot of course test all possible scalar field models, but we have done so for several potentials and obtained similar results. For brevity we will here only present the results for two scalar field potentials.
@@
$V=V_{-}e^{-\nu\left(1+\tanh{\left(\frac{\phi}{\sqrt{6\alpha}}\right)}\right)},$ (61)
@@
$\begin{split}\lambda&=\frac{\nu}{\sqrt{6\alpha}}\left(1-\tanh^{2}{\left(\frac{\phi}{\sqrt{6\alpha}}\right)}\right),\\ \lambda_{,\phi}&=-\frac{\nu}{3\alpha}\tanh{\left(\frac{\phi}{\sqrt{6\alpha}}\right)}\left(1-\tanh^{2}{\left(\frac{\phi}{\sqrt{6\alpha}}\right)}\right).\end{split}$ (62)
To obtain an explicit model, we set the model parameters $\alpha$ and $\nu$ as follows: $(\alpha,\nu)=(\frac{7}{3},128)$. We considered these models in [40] in the context quintessential inflation, where inflationary considerations determined $V_{-}$, see [50] and [51]. The inclusion of radiation yielded an integral, which resulted in that only one of the thawing solutions in the 1-parameter set characterized by $\phi_{*}$ was compatible with the chosen values of $V_{-}$, $H_{0}$, $\Omega_{\gamma 0}$ and $\Omega_{\mathrm{m}0}$. This solution was found by varying the frozen scalar field value, $\phi_{*}$, at the line of fixed points from which the thawing solutions originated, which resulted in $\phi_{*}=9.594$.
In the present case where we neglect radiation the integral does not exist and then $H_{0}$ and $\Omega_{\phi 0}=0.68$ do not impose any restrictions on $\phi_{*}$ at the line of fixed points $(\phi,\Sigma_{-},\Omega_{\phi})=(\phi_{*},0,0)$ from which the thawing quintessence oribits originate. However, due to that radiation is negligible when quintessence affects the evolution, the present approximations are excellent also when radiation is included as in [40]. For this reason we choose $\phi_{*}=9.594$. Inserting this value into (62) enables the subsequent computation of $\gamma$ and $\beta$ by means of (46) and (47), which results in $(\gamma,\beta)=(1.095,0.789)$, thereby making the approximation numerically explicit. To obtain the relative errors we numerically integrate the dynamical system (39) and the decoupled equation for $E_{\phi}$,

### v3

@@ 3.2 Tracking quintessence
Next, we introduce a dynamical system that is suitable for deriving the parameters $\gamma$ and $\beta$ for tracking quintessence. Tracking quintessence for asymptotically inverse power law potentials requires that in the vicinity of $\phi=0$ the potential and $\lambda$ are monotonically decreasing, i.e., when $\lambda(\phi)>0$ and $\Gamma(\phi)>1$, as follows from (38) and (48). We now introduce a system of equations that are based on the following new variables: [^13] The variables $\tilde{\phi}$ and $v$ are slightly different than the variables $\bar{\varphi}$ and $v$ in [37], which were adapted to a global state space setting, while the present ones are designed for fast convergence for the tracker orbit and the explicit appearance of $\Gamma$ in the equations and results.
@@
The dynamical system (52) admits $\tilde{\phi}=0$ (and thereby $\phi=0$) as an invariant boundary subset. As shown in [37], in regularizing variables, like the present ones, tracking quintessence is associated with a tracker orbit that originates from a tracker fixed point, $\mathrm{T}$, on the $\tilde{\phi}=0$ boundary, which in the present variables is given by
@@
Notably $\mathrm{T}$ is a hyperbolic saddle with the ‘tracker orbit’ as its unstable manifold and $\tilde{\phi}=0$ being its 2D stable manifold. This results in that an open set of nearby orbits is pushed toward the attractor orbit by the 2D invariant stable manifold boundary set $\tilde{\phi}=0$, leading to that the tracker orbit acts as an ‘attractor orbit’ that describes the evolution of the open attracted set of tracking/shadowing orbits during their quintessence evolution, see [37] for details.
To find the parameters $w_{\infty}$, $\gamma$ and $\beta$ we first note that the tracker fixed point value for $u=u_{\mathrm{T}}$ gives the past asymptotic tracker value for $w_{\phi}$ since
@@
$\Gamma(\tilde{\phi})\approx\Gamma^{(0)}+\Gamma^{(1)}\tilde{\phi}+\frac{1}{2}\Gamma^{(2)}\tilde{\phi}^{2}+\dots,\qquad\Gamma^{(n)}=\left.\frac{d^{n}\Gamma}{d\tilde{\phi}^{n}}\right|_{\tilde{\phi}=0},$ (56)
@@
where we recall that $\Gamma^{(0)}=1+p^{-1}$, where $p>0$ came from $V\propto\phi^{-p}$ when $\phi\rightarrow 0$, while $\Gamma^{(1)}$ and $\Gamma^{(2)}$ where defined in (56), $w_{\infty}=-2/(2+p)$, and where $\tilde{\phi}$ was defined in (50), i.e., $\tilde{\phi}=(V_{,\phi}/V)^{-2}$. In the expression for $k^{(2)}$ we assume that $k\neq 0$, since the special class of models with $k=0$, i.e. $\Gamma^{(1)}=\frac{3}{2}w_{\infty}$, needs special treatment; note, however, that this class include models with $w_{\phi}=\mathrm{constant}=w_{\infty}$, see [37] for details. [^14] Sahni and Starobinsky (2000) [50], equation (121), and Urena-Lopez and Matos (2000) [51], investigated models with the potential $V=V_{0}\sinh^{-p}(\nu\phi)$ and showed that these models admit a special solution (which turns out to be the tracker orbit) for which $w_{\phi}$ is constant, i.e., they yield $w$CDM spacetimes with $-1<w_{\phi}=w_{\infty}=w<0$.
Using (46) for tracking quintessence in connection with (30b) and (33) yields the following $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T)$ based master tracking quintessence approximations
$w_{\phi}$ $\approx[1/1]_{w_{\mathrm{DE}}}(T)=-\left(1-\frac{(\gamma-1)T}{1+\beta T}\right),$ (58a)
$\Omega_{\phi}$ $\approx\frac{\Omega_{\phi 0}}{\Omega_{\mathrm{m}0}e^{3w_{\infty}N}\left(\frac{1+\beta T}{1+\beta T_{0}}\right)^{\frac{\gamma-1}{\beta}}+\Omega_{\phi 0}},$ (58b)
$E_{\phi}^{2}$ $\approx\Omega_{\mathrm{m}0}e^{-3N}+\Omega_{\phi 0}e^{-3(1+w_{\infty})N}\left(\frac{1+\beta T_{0}}{1+\beta T}\right)^{\frac{\gamma-1}{\beta}},$ (58c)
where we recall that $T=T_{0}\exp(-3w_{\infty}N)$, $T_{0}=\Omega_{\phi 0}/(1-\gamma\Omega_{\phi 0})$ while $\gamma$ and $\beta$ were given in the previous eq. (57). Setting $\beta=\gamma-1$ in (58) in the above expressions yields the simpler, but less accurate, $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T)$ based tracking quintessence approximations. Again, to express our approximations in the redshift $z$, just use the simple relation $z=\exp(-N)-1$. Section 5 provides a specific tracking quintessence example, the inverse power law potential $V\propto\phi^{-p}$, $p>0$.
4 The Chevallier-Polarski-Linder parametrization
The so-called CPL parameters [27, 28] parametrize $w_{\mathrm{DE}}$ according to
$w_{\mathrm{DE}}=w_{\mathrm{CPL}}:=w_{0}+w_{a}\left(1-\frac{a}{a_{0}}\right),$ (59)
which can be viewed as a truncated Taylor expansion in $a/a_{0}$ at the present time (i.e. at $a/a_{0}=1$), where $a_{0}$ is usually set to one, which when expressed in the redshift $z$ takes the form $w_{\mathrm{CPL}}=w_{0}+w_{a}\left(\frac{z}{1+z}\right)$. Expressing the above equation in the $e$-fold time $N$ yields
$w_{\mathrm{CPL}}=w_{\infty}^{\mathrm{CPL}}-w_{a}e^{N},$ (60)
where $w_{\infty}^{\mathrm{CPL}}:=\lim_{N\rightarrow-\infty}w_{\mathrm{CPL}}$, from which it follows that
$w_{\infty}^{\mathrm{CPL}}=w_{0}+w_{a}.$ (61)
Inserting (60) into (7) yields
$f=e^{-3w_{\infty}^{\mathrm{CPL}}N}\cdot e^{-3w_{a}(1-e^{N})},$ (62)
which results in that (13) and (12) yield
$\Omega_{\mathrm{CPL}}$ $=\frac{\Omega_{\mathrm{DE}0}}{\Omega_{\mathrm{m}0}e^{3w_{\infty}^{\mathrm{CPL}}N}\cdot e^{3w_{a}(1-e^{N})}+\Omega_{\mathrm{DE}0}},$ (63a)
$E_{\mathrm{CPL}}^{2}$ $=\Omega_{\mathrm{m}0}e^{-3N}+\Omega_{\mathrm{DE}0}e^{-3(1+w_{\infty}^{\mathrm{CPL}})N}\cdot e^{-3w_{a}(1-e^{N})}.$ (63b)
We note that the expressions in (63) do not have simple single series expansion since they mix the functions $e^{N}$ and $e^{3w_{\infty}^{\mathrm{CPL}}N}$, which is in contrast to our expressions that are based on the single function $e^{-3w_{\infty}N}$, as is $w$CDM and $\Lambda$CDM when $w_{\infty}=-1$.
Usually the CPL parameters are cosmographically determined by observations, but sometimes the $w_{\mathrm{CPL}}(N)$ model with the parameters $w_{a}$ and $w_{0}$, where one of these parameters can be replaced with $w_{\infty}^{\mathrm{CPL}}$, is fitted to a numerically calculated solution of the quintessence field equations (due to that most scalar field potentials do not result in explicit solutions) with $w_{\phi}(N)$ at some time $N_{\mathrm{fit}}$ by requiring that
$w_{\mathrm{CPL}}^{\prime}(N_{\mathrm{fit}})$ $=-w_{a}\,e^{N_{\mathrm{fit}}}=w_{\phi}^{\prime}(N_{\mathrm{fit}}),$ (64a)
$w_{\mathrm{CPL}}(N_{\mathrm{fit}})$ $=w_{\infty}^{\mathrm{CPL}}-w_{a}\,e^{N_{\mathrm{fit}}}=w_{\phi}(N_{\mathrm{fit}}),$ (64b)
from which we can obtain $w_{a}$, $w_{0}$ and $w_{\infty}^{\mathrm{CPL}}$ at some fitting time $N_{\mathrm{fit}}$ as follows:
$w_{a}$ $=-e^{-N_{\mathrm{fit}}}\,w_{\phi}^{\prime}(N_{\mathrm{fit}}),$ (65a)
$w_{0}$ $=w_{\phi}(N_{\mathrm{fit}})-w_{\phi}^{\prime}(N_{\mathrm{fit}})(1-e^{-N_{\mathrm{fit}}}),$ (65b)
$w_{\infty}^{\mathrm{CPL}}$ $=w_{\phi}(N_{\mathrm{fit}})-w_{\phi}^{\prime}(N_{\mathrm{fit}}),$ (65c)
which if $N_{\mathrm{fit}}=0$ reduces to
$w_{a}$ $=-w_{\phi}^{\prime}(0),$ (66a)
$w_{0}$ $=w_{\phi}(0),$ (66b)
$w_{\infty}^{\mathrm{CPL}}$ $=w_{\phi}(0)-w_{\phi}^{\prime}(0).$ (66c)
In stark contrast to earlier work we will now use our quintessence approximations to analytically compute approximate observationally indistinguishable predictions for the CPL parameters. Using the for thawing and tracking quintessence unifying equations in (30) for $w_{\phi}(N_{\mathrm{fit}})\approx w_{\mathrm{DE}}(N_{\mathrm{fit}})$ result in rather complicated expressions. For brevity we therefore use (26) where we recall that $T=T_{0}\exp(-3w_{\infty}N)$, $T_{0}=\Omega_{\phi 0}/(1-\gamma\Omega_{\phi 0})$, where we have replaced $\Omega_{\mathrm{DE}0}$ with $\Omega_{\phi 0}$ since we now are concerned with quintessence. This leads to that the expression $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T)$ in (26b) results in
$w_{a}$ $=-3w_{\infty}\,e^{-N_{\mathrm{fit}}}\,X,$ (67a)
$w_{0}$ $=w_{\infty}-\left[1+3w_{\infty}(1-e^{-N_{\mathrm{fit}}})+\beta T_{\mathrm{fit}}\right]X,$ (67b)
$w_{\infty}^{\mathrm{CPL}}$ $=w_{\infty}-(1+3w_{\infty}+\beta T_{\mathrm{fit}})X,$ (67c)
where $X$ is defined as
$X=\frac{w_{\infty}(\gamma-1)T_{\mathrm{fit}}}{(1+\beta T_{\mathrm{fit}})^{2}},$ (68)
where $T_{\mathrm{fit}}=T_{0}\exp(-3w_{\infty}N_{\mathrm{fit}})$. As usual, the corresponding results for the $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T)$ are easily obtained from the above ones by replacing $\beta$ with $\gamma-1$. This leads to simpler expressions, but not sufficiently accurate for observational indistinguishability.
At the present time as fitting time, i.e. $N_{\mathrm{fit}}=0$, the equations in (67) yield
$w_{a}$ $\approx\frac{-3w_{\infty}^{2}(\gamma-1)(1-\gamma\Omega_{\phi 0})\Omega_{\phi 0}}{(1+(\beta-\gamma)\Omega_{\phi 0})^{2}},$ (69a)
$w_{0}$ $\approx w_{\infty}-\frac{w_{\infty}(\gamma-1)\Omega_{\phi 0}}{1+(\beta-\gamma)\Omega_{\phi 0}},$ (69b)
$w_{\infty}^{\mathrm{CPL}}$ $=w_{0}+w_{a},$ (69c)
which hence constitute our analytic quintessence predictions for the CPL parameters at the present time when our expressions for $\gamma$ and $\beta$ for quintessence are inserted. It follows that
$\frac{w_{a}}{w_{0}-w_{\infty}}\approx\frac{3w_{\infty}(1-\gamma\Omega_{\phi 0})}{1+(\beta-\gamma)\Omega_{\phi 0}}=3w_{\infty}-\frac{3w_{\infty}\beta\,\Omega_{\phi 0}}{1+(\beta-\gamma)\Omega_{\phi 0}},$ (69d)
for the $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T)$ case. As before, the expressions obtained from $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T)$ at $N_{\mathrm{fit}}=0$ are obtained by replacing $\beta$ with $\gamma-1$, where we note that $1+(\beta-\gamma)\Omega_{\phi 0}$ then reduces to $1-\Omega_{\phi 0}=\Omega_{\mathrm{m}0}$. This leads to the following simple approximate estimate (which, unfortunately, is not at all as good as that in (69d))
$\frac{w_{a}}{w_{0}-w_{\infty}}\approx\frac{3w_{\infty}(1-\gamma\Omega_{\phi 0})}{\Omega_{\mathrm{m}0}}=3w_{\infty}-3w_{\infty}(\gamma-1)\left(\frac{\Omega_{\phi 0}}{\Omega_{\mathrm{m}0}}\right),$ (70)
which thereby is linear in $\gamma$ and equal to $3w_{\infty}$ for $w$CDM since $\gamma=1$ in this case. Thawing quintessence, for which $w_{\infty}=-1$, yields
$\frac{w_{a}}{w_{0}+1}=-3+3(\gamma-1)\left(\frac{\Omega_{\phi 0}}{\Omega_{\mathrm{m}0}}\right)=-3+\frac{8}{9}\left(\frac{\Omega_{\phi 0}}{\Omega_{\mathrm{m}0}}\right)\epsilon_{*},$ (71)
due to (44). Thus if $\epsilon_{*}\approx 0$ then $w_{a}/(w_{0}+1)\approx-3$, however, viable thawing quintessence models typically admit a range of $\epsilon_{*}$ up to ${\cal O}(1)$, which results in a range for $w_{a}/(w_{0}+1)\in[-3,-1]$, cf. the discussion in [32], which requires the approximation in (69d) with $w_{\infty}=-1$ in order to achieve observational indistinguishability between a numerically computed thawing quintessence solution and our approximated one.
@@ 5 Numerical comparisons
In this section we calculate numerical relative errors for various types of approximations for quintessence, i.e.,
@@
where we for brevity consider the relative errors for $w_{\phi}(z)$, $\Omega_{\mathrm{m}}(z)=1-\Omega_{\phi}(z)$ and $E_{\phi}(z)$ (recall that $1+z=\exp(-N)$). Apart from calculating the relative errors of our approximations for thawing and tracking quintessence, we will also calculate relative errors for approximations found in the literature, described in appendix A; in addition, we calculate the relative errors of the CPL approximation for quintessence, described in section 4, for which the CPL parameters $w_{0}$ and $w_{a}$ must be calculated numerically. This in stark contrast to our purely analytic approximations which yield predictions for $w_{\phi}(z)$, $\Omega_{\mathrm{m}}(z)=1-\Omega_{\phi}(z)$, $E_{\phi}(z)$ and the CPL parameters $w_{0}$ and $w_{a}$. Furthermore, in order to illustrate that our most accurate approximations eliminate the need for numerical CPL quintessence calculations, we compute
$\begin{split}\Delta(F_{[0/1]})&=\left|\left|\frac{F_{CPL}-F_{num}}{F_{num}}\right|-\left|\frac{F_{CPL}-F_{[0/1]}}{F_{[0/1]}}\right|\right|,\\ \Delta(F_{[1/1]})&=\left|\left|\frac{F_{CPL}-F_{num}}{F_{num}}\right|-\left|\frac{F_{CPL}-F_{[1/1]}}{F_{[1/1]}}\right|\right|,\end{split}$ (73)
where $F=w_{\phi}(z)$, $F=\Omega_{\mathrm{m}}(z)=1-\Omega_{\phi}(z)$ and $F=E_{\phi}(z)$, while the subscripts [0/1] and [1/1] refer to our two approximations.
We cannot of course test all possible scalar field potentials, but we have done so with several potentials and obtained similar results. For brevity we will here only present the results for two specific scalar field potentials, one illustrating relative approximation errors for thawing quintessence and one for tracking quintessence.
@@
$V=V_{-}e^{-\nu(1+\bar{\phi})},\qquad\bar{\phi}=\tanh\left(\frac{\phi}{\sqrt{6\alpha}}\right),$ (74)
@@
$\lambda=\frac{\nu}{\sqrt{6\alpha}}\left(1-\bar{\phi}^{2}\right),\qquad\lambda_{,\phi}=-\left(\frac{\nu}{3\alpha}\right)\bar{\phi}\left(1-\bar{\phi}^{2}\right).$ (75)
As a consequence (42) yields
$\epsilon_{*}=\frac{\nu^{2}}{12\alpha}\left(1-\bar{\phi}_{*}^{2}\right)^{2},\qquad\eta_{*}=\frac{\nu}{6\alpha}\left(1-\bar{\phi}_{*}^{2}\right)\left[\nu\left(1-\bar{\phi}_{*}^{2}\right)+2\bar{\phi}_{*}\right],$ (76)
where $\bar{\phi}_{*}(\phi_{*})$, $\phi_{*}\in[-\infty,\infty]$, represents a freezing value of $\phi$ during the matter dominated epoch for the 1-parameter set of attractor thawing solutions with model parameters $\alpha$ and $\nu$ ($V_{-}$ represents an overall scaling parameter and do not appear in the dimensionless quantities $w_{\phi}(z)$, $\Omega_{\mathrm{m}}(z)=1-\Omega_{\phi}(z)$ and $E_{\phi}(z)$).
The parameters $\gamma$ and $\beta$ follow trivially from (47) and are given by
$\gamma=1+\frac{2\nu^{2}}{81\alpha}\left(1-\bar{\phi}_{*}^{2}\right)^{2},\qquad\beta=\frac{3}{5}+\frac{4\nu}{405\alpha}\left(1-\bar{\phi}_{*}^{2}\right)\left[2\nu\left(1-\bar{\phi}_{*}^{2}\right)+9\bar{\phi}_{*}\right],$ (77)
where (47) also yields our $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ based approximations while setting $\beta=\gamma-1$ into (47) first and then the above expression for $\gamma$ into the resulting expression yields our simpler but less accurate $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T(z))$ based approximations.
To compare with numerics we have to specify the model parameters $\alpha$ and $\nu$. We follow [38] and set them to $(\alpha,\nu)=(\frac{7}{3},128)$. We also need to specify which one of the 1-parameter set of thawing quintessence attractor solutions, parametrized by $\phi_{*}$, we consider. These models were discussed in [38] in the context quintessential inflation, where inflationary considerations determined the overall energy scale $V_{-}$, see [53] and [54]. The inclusion of radiation yielded an integral, which resulted in that only one of the thawing quintessence attractor solutions in the 1-parameter set characterized by $\phi_{*}$ was compatible with the chosen values of $V_{-}$, $H_{0}$, $\Omega_{\gamma 0}$ and $\Omega_{\mathrm{m}0}$. This solution was found by varying the frozen scalar field value, $\phi_{*}$, at the line of fixed points from which the thawing quintessence attractor solutions originated, which resulted in $\phi_{*}=9.594$. In the present case where we neglect radiation the integral does not exist and then $H_{0}$ and $\Omega_{\phi 0}=0.68$ do not impose any restrictions on $\phi_{*}$ at the line of fixed points $(\phi,\Sigma_{-},\Omega_{\phi})=(\phi_{*},0,0)$ from which the thawing quintessence attractor solutions originate. However, due to that radiation is negligible when quintessence affects the evolution, the present approximations for e.g. $w_{\phi}(z)$, $\Omega_{\phi}(z)$, $E_{\phi}(z)$ are excellent also when radiation is included as in [38]. For this reason, although not necessary, we choose $\phi_{*}=9.594$ in order to make contact with our previous work in [38]. Inserting $\phi_{*}=9.594$ and $(\alpha,\nu)=(\frac{7}{3},128)$ into (74) and (77) result in
$(\gamma,\beta)=(1.095,0.789),$ (78)
thereby making the approximation numerically explicit. To obtain the relative errors we numerically integrate the dynamical system (37) and the decoupled equation for $E_{\phi}=H_{\phi}/H_{0}$,
@@
The relative errors of $w_{\phi}(z)$, $\Omega_{\mathrm{m}}(z)=1-\Omega_{\phi}(z)$ and $E\phi(z)$ for several approximations are depicted in Figure 1. These include analytic approximations such as our $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T(z))$ and $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ based approximations, as well as the approximations by Scherrer and Sen [55] and Chiba [56], given by the equations (88) and (94), respectively, in appendix A. In addition we give the CPL approximation based on the numerically calculated values for $w_{0}$ and $w_{a}$ at $z=0$, which thereby, of course, results in no error for $w_{\phi}$ at $z=0$.
As can be seen, the earlier analytic approximations by Scherrer and Sen [55] and Chiba [56] [^15] Here we have treated these approximations in the same way as our approximations, i.e. as analytic predictive approximations. This is not how they were dealt with in the original references. There numerical calculations picked out a particular solution in the 1-parameter set of thawing solutions, which results in less errors at the price of loosing predictive power. and the more simple and consistent $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T(z))$ approximant based approximations have similar relative errors for the physical observables $\Omega_{\mathrm{m}}$ and $E_{\phi}$, with respective $\sim 3\%$ and $\sim 1.5\%$ maxima, while the $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ approximant yields a maximum relative error of $\sim 0.1\%$ for $\Omega_{\mathrm{m}}$ and an even smaller error for $E_{\phi}$, making this approximation observationally indistinguishable from numerical results. Not even the numerically based CPL approximation, with no predictive power at $z=0$, is as good as our analytic $w_{\phi}\approx[1/1]_{w_{\mathrm{DE}}}(T(z))$ based approximations.
Note that if we instead of using $T_{0}=\Omega_{\phi 0}/(1-\gamma\Omega_{\phi 0})$ use the numerically computed $w_{0}=w_{\phi 0}$ in $T_{0}$ as described in eq. (27) we get even more accurate results than our previous approximations and the CPL based approximations, but then this results in that the purely analytic predictions of our quintessence approximations are lost.